OpenStudy (anonymous):

Hector needs to borrow $20000 to pay a contractor for a home addition. The bank is offering him a 4.8% simple interest loan and is charging $4800 in interest during the life of the loan. How long is the life of the loan?

OpenStudy (bahrom7893):

What is the formula for simple interest?

OpenStudy (anonymous):

I=prt

OpenStudy (anonymous):

Is this correct ?

OpenStudy (bahrom7893):

Where did you get the equation from?

OpenStudy (anonymous):

My algerbra book

OpenStudy (bahrom7893):

Okay, then we can assume it is correct. So now, what do l, p, r, and t stand for?

OpenStudy (anonymous):

whoops typo lol algebra bok my bad lol I stands for intrest T stands for time R stands for Rate and P stand for principal so basically it's this p = Principal, or the money in your savings account r = The rate of interest, usually in a % t = Time. The time between interest payments.

OpenStudy (anonymous):

I need to find T or time how exactly do I do that?

OpenStudy (bahrom7893):

So what do you know right now? You know Interest, Rate of Interest, and Principal

OpenStudy (bahrom7893):

Plug them into your formula

OpenStudy (anonymous):

principal is 20,00 and the rate 4.8% which is 4800 whe converted I did this p/r which was 20,00/4800 and got 4.16666667so I'm guessing the answer for T is 4yrs correct?

OpenStudy (anonymous):

Am I right or wrong ?

OpenStudy (whpalmer4):

4.8% is not 4800 when converted \[4.8\% = 4.8\% * 1/100\% = 0.048\]

OpenStudy (whpalmer4):

What will be the interest on $1,000 for 1 year?

OpenStudy (anonymous):

1000/0.048=2.08333333

OpenStudy (anonymous):

so it's 2 yrs?

OpenStudy (whpalmer4):

No, no.... If you are paying me 10% interest on a loan for 1 period, you pay me 10% of the principal as interest. Why are you dividing by the interest rate? You need to multiply!

OpenStudy (whpalmer4):

interest = principal * rate * time

OpenStudy (anonymous):

Oh so that had nothing to do with my math question you were just give me and example. My bad sorry XD The formula is P=RT correct so if I divide P/R won't that give me T . I see it seems I've skipped a letter

OpenStudy (anonymous):

so it's I=PRT my bad lol XD

OpenStudy (anonymous):

I don't get why do have to multiply I'm trying to find time not intrest I already have that so shouldn't I divide something to get time? I'm so confused right now I feel like I complete dunce right now sigh T-T

OpenStudy (whpalmer4):

You need to understand how interest works...you are paying a fraction of the value of the principal each period. That fraction is the interest rate (expressed as a decimal) times the principal. If there are multiple periods, you also multiply by the number of periods. So, the interest on your loan will be 4.8% * $20,000 * T where T is the number of years. You know that the total amount of interest paid is $4,800. $4,800 = 4.8% * $20,000 * T 4800 = 0.048 * 20000 * T Can you find the value of T?

OpenStudy (johnweldon1993):

I think what YOU want to do....is rearrange the equation for time You know that I = PRT (and yes this is correct) So to find time....you simply divide both sides of the equation by PR to isolate T I = PRT ------- PR PR Now you have \[T = \frac{ I }{ PR }\] You know I = 4800 P = 20000 and R = .048 So plugged into your equation you have \[T = \frac{ 4800 }{ 20000 * .048 }\] What does T equal??